OpenStudy (anonymous):

locate the absolute extrema of the function on the closed interval f(x) = 2x/(x^2+1), [-2,2] if someone could help with the process, it'd be great since I may have gone somewhere wrong

OpenStudy (anonymous):

take the derivative, set it equal to zero and solve then check the zeros of the derivative, also the endpoints of the interval largest is max, smallest is min

OpenStudy (anonymous):

yeah I understand that but I got a really weird answer when I tried to find the derivative

OpenStudy (anonymous):

I used the quotient rule and got x = sqrt (-1/3)

OpenStudy (anonymous):

after making the derivative equal to 0

OpenStudy (anonymous):

i think your derivative is wrong problem was cooked so that it is nice and easy to find the zeros

OpenStudy (anonymous):

\[f'(x)=\frac{-2(x^2-1)}{(x^2+1)^2}\] zeros are easy and easy to check too i think

OpenStudy (anonymous):

the derivative that I got before trying to find the 0 value was (6x^2+2)/(X^2+1)^2

OpenStudy (anonymous):

I know I'm asking for a lot but is there any way that you could guide me through the process so that I can see what I did wrong?

OpenStudy (anonymous):

numerator should be \[2(x^2+1)-2x(2x)\]

OpenStudy (anonymous):

yep thats what I got in the numerator

OpenStudy (anonymous):

multiply out and get \[2x^2+2-4x^2=-2x^2+2=-2(x^2-1)\]

OpenStudy (anonymous):

so zeros are \(-1,1\) both of which are in the interval

OpenStudy (anonymous):

i am guessing you made an algebra mistake between \[2(x^2+1)-2x(2x)\]and \[-2(x^2-1)\]

OpenStudy (anonymous):

yeah you're definitely correct, instead of subtracting the two I added, I got it mixed up with the product rule

OpenStudy (anonymous):

I understand this fully thanks for taking the time to explain it to me :)